OpenStudy (18jonea):

Is the sequence –9.9, –13, –16, –19.2, ... arithmetic? If so, identify the common difference. yes, –3.2 yes, –2.9 yes, –3.1 no I think it is c

OpenStudy (18jonea):

@sweetburger

OpenStudy (18jonea):

@surjithayer

OpenStudy (18jonea):

?

OpenStudy (sweetburger):

I don't think c is correct as the 3rd number of the sequence would have to be -16.1

OpenStudy (anonymous):

-13-(-9.9)=-13+9.9=-3.1 -16-(-13)=-16+13=-3 -19.2-(-16)=-19.2+16=-3.2 now you can judge whether it is an arithmetic sequence or not.

OpenStudy (mathmale):

Please determine what the difference is between -9.9 and -13, -13 and -16, and so on. If the difference is not the same from one to the next, this is not an arith sequence.

OpenStudy (18jonea):

So it is D @surjithayer

OpenStudy (anonymous):

correct.

OpenStudy (18jonea):

Ok I have one more that i think I have the answer for could you please check @surjithayer

OpenStudy (anonymous):

sure i will try.

OpenStudy (18jonea):

Write an explicit and a recursive formula for the sequence –7, –5.6, –4.2, –2.8, –1.4, .... explicit: an = –8.4 + 1.4n; recursive: an = an – 1 + 1.4, where a1 = –7 explicit: an = –8.4 + 1.4n; recursive: an = an – 1 + 1.4n, where a1 = –7 explicit: an = –7 + 1.4n; recursive: an = an – 1 + 1.4n, where a1 = –7 explicit: an = –7 + 1.4n; recursive: an = an – 1 + 1.4, where a1 = –7 I think it is a

OpenStudy (18jonea):

@surjithayer

OpenStudy (anonymous):

correct.

OpenStudy (18jonea):

Thank You!

OpenStudy (anonymous):

yw